OpenStudy (anonymous):

A printer finds for every 12 pages of newspaper produced, there are 3 misspellings. If the printer produces 13,200 pages of newspaper this week, what is the best prediction for how many misspellings there will be?

OpenStudy (rituraj):

12 pages - 3 misspellings 1 page - 3/12 = 1/4 misspellings So 13200 pages - 13200/4 = 3300 misspellings